A child pushes a 31-kg shopping cart by a force of 120 N in a supermarket at angle of 25 degrees above the horizontal. If the cart experiences a friction force of 50 N that acts simultaneously on it, calculate the magnitude of its horizontal acceleration.
Bonus: Calculate the normal force acting on the cart.
1
Expert's answer
2020-03-04T10:10:19-0500
"ma=F\cos{25}-F_f"

"31a=120\cos{25}-50"

"a=1.9\frac{m}{s^2}"

The normal force acting on the cart:


"N=mg-F\sin{25}"

"N=(31)(9.8)-120\sin{25}=250\ N"